OpenStudy (anonymous):

plz help urgent What is the sum of the geometric sequence 8, –16, 32 … if there are 15 terms?

OpenStudy (anonymous):

@JA1

OpenStudy (anonymous):

@Jasmineflvs

OpenStudy (texaschic101):

what is the common difference ? Can you answer that ?

OpenStudy (anonymous):

8

OpenStudy (texaschic101):

no...this is geometric not arithmetic.....what does 8 multiply by to = -16, and what does -16 multiply by to equal 32 ?

OpenStudy (anonymous):

oh ok that would be -2

OpenStudy (anonymous):

i dont understand on how to solve it i have reveowed all of my notes but i am completely lost

OpenStudy (anonymous):

i think i just figured it out do i just keep multiplying by -2 until i get to the 15th term?

OpenStudy (texaschic101):

correct. Now we will use the sum formula... sn = a1 (1 - r^n) / (1 - r) n = 15 r = -2 a1 = 8 s15 = 8(1 - (-2^15) / (1 - (-2) s15 = 8 (1 + 2^15) / 3 s15 = 8(32769)/3 s15 = 262152/3 s15 = 87384

OpenStudy (anonymous):

oh ok i see what to do now my teacher wont help me i apperciate your help and patience with me.

OpenStudy (texaschic101):

you could just keep multiplying by -2 and when you get to the 15th term add everything that you have...which would take awhile

OpenStudy (texaschic101):

This might help you... http://www.regentsprep.org/Regents/math/algtrig/ATP2/GeoSeq.htm

OpenStudy (texaschic101):

just remember, when it is geometric, the common difference is gonna be result of multiplying, but in an arithmetic sequence, the common difference will be result of adding

OpenStudy (texaschic101):

any more questions ?

OpenStudy (anonymous):

not right at this moment i dont but if i do i will definitrly ask you.@texaschic101

OpenStudy (texaschic101):

I will be happy to help, if I can :)